### v1

# Bidirectional Prototype-Reward co-Evolution for Test-Time Adaptation of Vision-Language Models

Abstract
Test-time adaptation (TTA) is crucial in maintaining Vision Language Models(VLMs)’ performance when facing real-world distribution shifts, particularly when the source data or target labels are inaccessible. Existing TTA methods rely on CLIP’s output probability distribution for feature evaluation, which can introduce biases under domain shifts. This misalignment may cause features to be misclassified due to text priors or incorrect textual associations. To address these limitations, we propose Bidirectional Prototype-Reward co-Evolution (BPRE), a novel TTA framework for VLMs that integrates feature quality assessment with prototype evolution through a synergistic feedback loop. BPRE first employs a Multi-Dimensional Quality-Aware Reward Module to evaluate feature quality and guide prototype refinement precisely. The continuous refinement of prototype quality through Prototype-Reward Interactive Evolution will subsequently enhance the computation of more robust Multi-Dimensional Quality-Aware Reward Scores. Through the bidirectional interaction, the precision of rewards and the evolution of prototypes mutually reinforce each other, forming a self-evolving cycle. Extensive experiments are conducted across 15 diverse recognition datasets encompassing natural distribution shifts and cross-dataset generalization scenarios. Results demonstrate that BPRE consistently achieves superior average performance compared to state-of-the-art methods across different model architectures, such as ResNet-50 and ViT-B/16. By emphasizing comprehensive feature evaluation and bidirectional knowledge refinement, BPRE advances VLM generalization capabilities, offering a new perspective on TTA.
1 Introduction
Figure 1: A comparative analysis of cross-dataset generalization performance between CLIP models pre-trained with ResNet-50 (a) and ViT-B/16 (b) image encoders. BPRE demonstrates competitive results across ten diverse datasets, showcasing its robust generalization capabilities. (c) Bidirectional Prototype-Reward co-Evolution diagram.
Recent advances in Vision-Language Models (VLMs) [2, 38, 18] have opened new opportunities to integrate human language into various computer vision tasks. Despite the unprecedented zero-shot generalization capability achieved by VLMs such as CLIP [23], existing VLMs suffer from a large performance drop when encountering domain shifts, which is due to the inconsistent data distribution between the source domain and the target domain. However, in real-world applications, the inevitable domain discrepancies caused by dynamic and ever-changing environments make domain adaptation of VLMs an urgent problem.
To address this issue, some researchers have developed efficient adaptation methods [40, 45, 3] that fine-tune the VLMs using limited target domain samples. However, the requirement for target domain annotations remains a critical bottleneck, especially in dynamic environments with frequent distribution shifts [24, 36]. Recently, several studies have explored the test-time adaptation (TTA) for VLMs, aiming to adaptively optimize the model at a low cost during the test phase. Shu et al. [25] introduced test-time prompt tuning (TPT) that optimizes prompts for individual test samples to enhance out-of-distribution generalization. This framework was enhanced by DiffTPT [8] through diffusion-based augmentations, while Karmanov et al. [14] proposed a training-free approach using dynamic visual caches. More recently, Zhang et al. [37] developed a dual prototype evolving strategy that accumulates task-specific knowledge from both visual and textual modalities during testing.
Overall, most of the existing TTA works [40, 14, 37] have highlighted the importance of utilizing historical feature information to accumulate task-specific knowledge for each class by Caches and Prototypes. This approach aims to effectively construct accurate and discriminative representations of target classes, which are crucial for enhancing the model’s robustness to distribution shifts, as well as improving its efficiency and performance during the test-time phase. Current methods [14, 37] primarily rely on CLIP’s output probability distribution, derived from the fusion of Text Features and Image Features, to evaluate feature quality, inherently using the text modality as a reference. However, as shown in Fig. 2, distribution shifts during testing may cause the text encoder to misalign with the semantic space of the image modality, leading to biases in feature selection and evaluation. On one hand, text modality priors cause the model to favor features that strongly align with pre-trained textual descriptions, even when they lack sufficient visual information, while visually rich samples with weaker textual representations may be misclassified (Fig. 2(a)). On the other hand, misalignment of textual descriptions can introduce feature evaluation biases, where CLIP may incorrectly associate generalized textual descriptions (e.g.,“picture”) with certain image features, leading the model to mistakenly consider them high-quality samples while overlooking their true semantics, thereby affecting feature selection accuracy (Fig. 2(b)).
Based on these insights, we propose a novel test-time VLM adaptation framework: Bidirectional Prototype-Reward co-Evolution (BPRE), which iteratively refines task-specific knowledge, as illustrated in Fig. 3. Unlike existing approaches, BPRE introduces the Multi-Dimensional Quality-Aware Reward Module, an evaluation method that integrates intrinsic Image Feature properties. By leveraging visual information while incorporating Semantic Similarity and Feature Diversity as key criteria, this module provides a more comprehensive assessment of Image Feature quality, reducing excessive dependence on the text modality. Semantic Similarity ensures selected features align with the task’s semantic space, mitigating errors from textual biases, while Feature Diversity enhances coverage across varied data distributions, preventing overfitting and improving prototype selection during test-time adaptation. Additionally, the Prototype-Reward Interactive Evolution module adaptively updates historical features and Prototypes using Reward scores. This bidirectional reinforcement loop allows refined prototypes to enhance reward estimation, forming a self-reinforcing optimization process that boosts robustness against noise and facilitates interpretable cross-modal learning, yielding more discriminative and stable representations.
Figure 2: (a) Dog/cat images, despite lacking visual details, may still be prioritized due to strong textual representations, while visually rich samples like a window may be misclassified as the nearest known category, such as a door, due to weak textual alignment. (b) Certain image regions that are visually associated with the term ”picture” may be erroneously emphasized during feature evaluation.
The test-time generalization of BPRE is evaluated across 15 diverse recognition datasets under natural distribution shifts and cross-dataset generalization. As shown in Fig. 1, BPRE demonstrates competitive performance in cross-dataset generalization with both ResNet-50 and ViT-B/16 architectures.
In summary, our contributions can be listed as:
  • We introduce the Multi-Dimensional Quality-Aware Reward Module, integrating Similarity, Confidence, and Diversity Reward Scores for comprehensive feature evaluation, improving prototype selection.
  • We propose the Prototype-Reward Interactive Evolution Module, establishing a bidirectional reinforcement mechanism where reward scores guide prototype evolution, and refined prototypes enhance reward estimation, improving model robustness.
  • Experiments across 15 diverse image recognition datasets confirm the competitive generalization performance of BPRE.
Figure 3: Overview of our proposed BPRE framework. Given text and image inputs, encoders $E_{t}(\cdot)$ and $E_{v}(\cdot)$ extract text features $f_{t}$ and visual features $f_{v}$. A probability distribution $\mathbb{P}_{CLIP}(y=y_{c}|X_{\mathtt{test}})$ is computed based on these features. A reward module then integrates similarity, confidence, and diversity metrics to derive a composite reward score $R_{final}$, guiding prototype updates. Finally, a residual-based refinement step adjusts the logits by compensating for the residuals between updated prototypes and test samples.
2 Related Works
2.1 Vision-Language Models
Large-scale VLMs such as CLIP [23] and SigLIP [35] have demonstrated remarkable success in visual understanding tasks by leveraging massive image-text pre-training [6, 39]. Their strong zero-shot transferability makes them effective backbones for tasks like image recognition [10, 19] and object detection [31, 17]. To better adapt VLMs to downstream tasks, researchers have proposed two main categories of methods: prompt learning methods [43, 42, 15, 45, 3] and adapter-based methods [40, 34, 9, 44]. Prompt learning methods like CoOp [43] and CoCoOp [42] focus on learning continuous prompt vectors with few-shot labeled data to bridge the domain gap between pre-training and downstream tasks. These methods achieve impressive performance but require labeled training samples that are often unavailable in real-world scenarios. Despite the success of adapter-based methods like Tip-Adapter [40] and TaskRes [34] in enhancing visual-textual representations, their dependence on labeled target data remains a critical bottleneck. Here we present a test-time adaptation framework that overcomes this limitation by enabling real-time model adaptation without prior exposure to target domain labels or training samples. Our approach facilitates robust model deployment across diverse, unfamiliar environments where labeled data collection is unfeasible.
2.2 Test-Time Adaptation
Test-time adaptation addresses domain shift challenges by enabling dynamic model adaptation during deployment using unlabeled test samples [28, 14, 37]. This approach has proven effective across various computer vision tasks, enhancing model robustness and maintaining performance when encountering out-of-distribution scenarios in real-world applications. Test-time adaptation of VLMs has recently gained significant research attention. TPT [25] pioneered this direction by leveraging consistency across augmented views of test samples. The following works enhanced this approach: DiffTPT [8] introduced diffusion-based augmentations, while C-TPT [33] focused on addressing calibration errors. Moving beyond single-sample adaptation, TDA [14] and DMN [41] introduced memory-based mechanisms to leverage information across multiple test samples, marking a shift from instance-level to collective adaptation strategies. DPE [37] proposes a dual-modality prototype evolution mechanism, which achieves more accurate feature modeling through dynamic updating and optimization of dual-modality feature representations during testing. However, existing methods including DPE [37] focus solely on prototype evolution, overlooking the mutual interaction between prototypes and sample quality. We address this by proposing a Bidirectional Prototype-Reward co-Envolution framework that enables quality-aware prototype evolution while ensuring consistent feature representation through mutual guidance.
3 Method
As shown in Fig. 3, we propose the BPRE framework to enhance CLIP’s zero-shot generalization at test time. Unlike prior methods [40, 14, 37], BPRE introduces a Bidirectional Prototype-Reward co-Evolution mechanism, where prototypes and rewards mutually refine each other. This enables quality-aware adaptation and consistent feature representation, dynamically updating prototypes with unlabeled test data $\mathcal{D}_{\mathtt{test}}$ without supervision.
3.1 Preliminary
Zero-Shot CLIP. CLIP [23] consists of dual encoders: a visual encoder $\mathcal{E}_{v}(\cdot)$ and a text encoder $\mathcal{E}_{t}(\cdot)$, which project images and text into a shared $d$-dimensional embedding space. CLIP performs zero-shot prediction for a $C$-class classification task by measuring similarities between image and text embeddings. Given a test image $X_{\mathtt{test}}\in\mathcal{D}_{\mathtt{test}}$ and class-specific text descriptions ${\mathcal{T}_{c}}$, the prediction process can be formulated as:
$f_{v}=\mathcal{E}_{v}(X_{\mathtt{test}}),\quad f_{t_{c}}=\mathcal{E}_{t}(\mathcal{T}_{c})$
$\mathbb{P}_{CLIP}(y=y_{c}|X_{\mathtt{test}})=\frac{\exp(\mathrm{sim}(f_{t_{c}},f_{v})/\tau)}{\sum_{t^{\prime}}\exp(\mathrm{sim}(f_{t^{\prime}},f_{v})/\tau)},$ (1)
where $\mathrm{sim}(\cdot,\cdot)$ denotes cosine similarity and $\tau$ is the temperature parameter.
Test-Time Prompt Tuning. TPT [25] improves CLIP’s zero-shot generalization by learning adaptive prompts from test samples. For each test image $X_{\mathtt{test}}$, TPT generates $N$ augmented views ${\mathcal{A}_{n}(X_{\mathtt{test}})}_{n=1}^{N}$ and aggregates their predictions based on confidence. Specifically, TPT selects predictions with entropy below threshold $t$ and computes their average within the top $\rho$-percentile:
$\mathbb{P}_{\mathtt{TPT}}(X_{\mathtt{test}})=\frac{1}{\rho N}\sum_{n=1}^{N}$  $\,\mathbf{1}\Big{[}\mathcal{H}\big{(}\mathbb{P}_{\mathtt{CLIP}}(\mathcal{A}_{n}(X_{\mathtt{test}}))\big{)}\leq t\Big{]}$
×$\quad\times\mathbb{P}_{\mathtt{CLIP}}(\mathcal{A}_{n}(X_{\mathtt{test}})),$ (2)
where $\mathcal{H}(p)=-\sum_{i=1}^{C}p_{i}\log p_{i}$ is the prediction entropy. The prompt is optimized to minimize the entropy of the final prediction: $\min\mathcal{H}(\mathbb{P}_{\mathtt{TPT}}(X_{\mathtt{test}}))$.
3.2 Multi-Dimensional Quality-Aware Reward
To effectively assess samples’ contributions to model performance and guide feature learning, we introduce a Multi-Dimensional Quality-Aware Reward Module, which dynamically assigns quality scores by integrating multiple characteristics.
Similarity Reward Score. In cross-modal learning, semantic consistency between sample features and category prototypes serves as a crucial indicator for sample quality evaluation. To this end, we design a similarity-based reward $R_{sim}$ that measures the alignment between sample features and prototype representations:
$R_{sim}=\frac{1}{|\mathcal{P}|}\sum_{p_{k}\in\mathcal{P}}\cos(\frac{f_{i}}{\|f_{i}\|_{2}},\frac{p_{k}}{\|p_{k}\|_{2}}),$ (3)
where $f_{i}\in\mathbb{R}^{d}$ is the feature vector of sample $x_{i}$, and $\mathcal{P}=\{p_{1},p_{2},\dots,p_{K}\}$ denotes the set of category prototypes, with each $p_{k}\in\mathbb{R}^{d}$ representing the prototype of category $k$. The function $\cos(\cdot,\cdot)$ computes the cosine similarity between the normalized feature vector $f_{i}$ and prototype $p_{k}$, while $\|\cdot\|_{2}$ denotes the $L_{2}$-norm for normalization. This similarity reward enhances feature discriminability by enforcing semantic consistency and aligning prototype evolution with category semantics.
Confidence Reward Score. To assess the model’s prediction confidence, we define a confidence reward $R_{conf}$ based on the entropy of the predicted probability distribution. Utilizing temperature-scaled normalization, $R_{conf}$ is derived from CLIP’s predicted probability distribution $\mathbf{P}_{CLIP}$:
$R_{conf}=1-\frac{-\sum_{i=1}^{C}\mathbb{P}_{CLIP}(X_{test})\log\mathbb{P}_{CLIP}(X_{test})}{\log C},$ (4)
where $C$ is the number of categories. This entropy-based confidence reward prioritizes high-certainty predictions, enhancing model reliability while fostering more robust feature representations.
Diversity Reward Score. To prevent representation collapse and encourage diverse feature exploration, we introduce a diversity reward $R_{div}$ based on historical feature comparisons. We maintain a fixed-size feature memory queue =$\mathcal{H}=\{h_{1},h_{2},\dots,h_{M}\}$, where each $h_{j}\in\mathbb{R}^{d}$ represents the average feature of a past batch, and $M$ is the queue size. For a given sample feature $f_{i}$, its novelty is measured by the maximum cosine similarity with historical features:
$S_{max}=\max_{h_{j}\in\mathcal{H}}\cos(\frac{f_{i}}{\|f_{i}\|_{2}},\frac{h_{j}}{\|h_{j}\|_{2}}),$ (5)
The diversity reward is then defined as:
$R_{div}=1-S_{max}.$ (6)
This mechanism enhances feature diversity while maintaining discriminability, improving the model’s generalization across downstream tasks.
By integrating the proposed reward components, we define an adaptive comprehensive reward function that evaluates sample quality while guiding prototype evolution:
$R=\lambda_{sim}R_{sim}+\lambda_{conf}R_{conf}+\lambda_{div}R_{div},$ (7)
where $\lambda_{sim}$, $\lambda_{conf}$, and $\lambda_{div}$ are hyperparameters that balance the contributions of similarity reward $R_{sim}$, confidence reward $R_{conf}$, and diversity reward $R_{div}$, respectively.
$R_{final}=R_{min}+(R-R_{min})*\eta_{\text{warm}},$ (8)
where $R_{min}$ is the minimum reward threshold, and $\eta_{\text{warm}}$ is a warmup factor. Initially, when $\eta_{\text{warm}}$ is small, $R_{final}$ remains close to $R_{min}$, stabilizing early adaptation. As $\eta_{\text{warm}}$ increases, $R_{final}$ gradually approaches $R$, enhancing adaptation while mitigating fluctuations.
This framework employs a comprehensive reward mechanism to optimize cross-modal transfer learning. By integrating similarity, confidence, and diversity metrics, it dynamically evaluates feature representations. The adaptive reward system, guided by a warmup phase and historical feature tracking, ensures stable learning signals while facilitating prototype evolution, effectively bridging reward estimation and prototype learning.
3.3 Prototype-Reward Interactive Evolution
Building on the Multi-Dimensional Quality-Aware Reward Module, we propose a dynamic prototype evolution mechanism that utilizes reward-guided learning to refine category-specific prototypes. By weighting feature contributions based on their computed rewards, this approach adaptively enhances prototype representations, fostering a more robust and discriminative feature space for cross-modal recognition.
Prototype Representation and Initialization. For a classification task with $C$ categories, we maintain a set of prototype representations in a $d$-dimensional image feature space: $v_{c}\in\mathbb{R}^{d}$. These prototypes act as learnable category-level representations, capturing essential class characteristics.
Figure 4: (a-d) t-SNE [29] visualizations of image features in the priority queues demonstrate that as additional samples are incorporated, class-specific features become more tightly clustered on the Food101 [1] dataset. (e) As the number of input samples grows, the reward score computed by the Multi-Dimensional Quality-Aware Reward Module increases.
At the start of training, prototypes are initialized using category-specific statistics. The visual prototype $v_{c}^{(0)}$ for category $c$ is set as the mean feature representation of all samples in that category:
$v_{c}^{(0)}=\frac{1}{|\mathcal{D}_{c}|}\sum_{x_{i}\in\mathcal{D}_{c}}f_{v},$ (9)
where $\mathcal{D}_{c}$ denotes the set of samples belonging to category $c$.
Quality-Aware Visual Prototype Evolution. Leveraging the comprehensive quality assessment derived from our multi-dimensional reward mechanism, we propose a quality-aware momentum update strategy for dynamic prototype evolution. This mechanism strategically incorporates sample contributions by utilizing their reward scores $R$ as quality indicators, thereby ensuring that samples with higher reward values exert greater influence on the prototype evolution process. Specifically, for category $c$, the visual prototype update at time step $t$ is formulated as:
$v_{c}^{(t+1)}=m\cdot v_{c}^{(t)}+(1-m)\cdot\sum_{x_{i}\in\mathcal{D}_{c}}w_{i}^{(t)}\cdot f_{v},$ (10)
where $m\in[0,1]$ is a momentum coefficient that stabilizes the update process by controlling the balance between historical information and new observations. The adaptive weight $w_{i}^{(t)}$ for each sample is computed through temperature-scaled softmax normalization:
$w_{i}^{(t)}=\frac{\exp(R_{final}/\tau)}{\sum_{j\in\mathcal{D}_{c}}\exp(R_{final}/\tau)}.$ (11)
where $R_{i}$ is the comprehensive quality score derived from our multi-dimensional reward mechanism, and $\tau$ is a temperature parameter that controls the sharpness of the weight distribution.
Method ImageNet ImageNet-A ImageNet-V2 ImageNet-R ImageNet-S Average OOD Average
CLIP-ResNet-50 [23] 58.16 21.83 51.41 56.15 33.37 44.18 40.69
Ensemble 59.81 23.24 52.91 60.72 35.48 46.43 43.09
CoOp [43] 63.33 23.06 55.40 56.60 34.67 46.61 42.43
TPT [25] 60.74 26.67 54.70 59.11 35.09 47.26 43.89
DiffTPT [8] 60.80 31.06 55.80 58.80 37.10 48.71 45.69
TDA [14] 61.35 30.29 55.54 62.58 38.12 49.58 46.63
TPS [27] 61.47 30.48 54.96 62.87 37.14 49.38 46.36
DMN-ZS [41] 63.87 28.57 56.12 61.44 39.84 49.97 46.49
DPE [37] 63.41 30.15 56.72 63.72 40.03 50.81 47.66
BPRE (Ours) 63.77 30.77 56.97 63.94 39.91 51.07 47.90
$\Delta$0.11 -0.29 +0.25 +0.22 -0.12 +0.26 +0.24
CLIP-ViT-B/16 [23] 66.73 47.87 60.86 73.98 46.09 59.11 57.20
Ensemble 68.34 49.89 61.88 77.65 48.24 61.20 59.42
CoOp [43] 71.51 49.71 64.20 75.21 47.99 61.72 59.28
TPT [25] 68.98 54.77 63.45 77.06 47.94 62.44 60.81
DiffTPT [8] 70.30 55.68 65.10 75.00 46.80 62.28 60.52
TDA [14] 69.51 60.11 64.67 80.24 50.54 65.01 63.89
TPS [27] 70.19 60.08 64.73 80.27 49.95 65.04 63.76
DMN-ZS [41] 72.25 58.28 65.17 78.55 53.20 65.49 63.80
DPE [37] 71.91 59.63 65.44 80.40 52.26 65.93 64.43
BPRE (Ours) 72.22 60.67 65.50 80.68 53.45 66.50 65.08
$\Delta$0.03 +0.56 +0.07 +0.28 +0.25 +0.57 +0.65
Table 1: Robustness Evaluation under Natural Distribution Shifts. We report comparative top-1 accuracy (%) across methods, utilizing CLIP’s ResNet-50 and ViT-B/16 architectures as visual encoders. Bold values indicate optimal performance.
Overall, the dual promotion mechanism creates a dynamic feedback loop between reward estimation and prototype evolution, facilitating their mutual enhancement. The mechanism operates through two complementary processes: (1) high-fidelity prototypes serve as anchor points for precise reward assessment, and (2) refined reward signals guide targeted prototype updates. As illustrated in Fig. 4, increasing the number of iterative samples results in tighter clustering of class-specific features, with a corresponding rise in reward scores from the Multi-Dimensional Quality-Aware Reward Module. This iterative interaction progressively enhances the framework’s ability to model fine-grained visual-semantic relationships and adapt to intra-class variations. By co-evolving prototypes and rewards, the mechanism improves robustness against noisy data and fosters interpretable cross-modal feature learning, ultimately producing more discriminative and generalizable representations for cross-modal recognition.
4 Experiments
We assess the BPRE framework through a thorough experimental protocol, focusing on three key dimensions: distributional robustness, cross-dataset generalization across 15 benchmarks, and inference efficiency. Additionally, systematic ablation studies are conducted to validate the core design principles guiding our architectural decisions.
Method Aircraft Caltech Cars DTD EuroSAT Flower Food101 Pets SUN397 UCF101 Average
CLIP-ResNet50 [23] 15.66 85.88 55.70 40.37 23.69 61.75 73.97 83.57 58.80 58.84 55.82
Ensemble 16.11 87.26 55.89 40.37 25.79 62.77 74.82 82.97 60.85 59.48 56.63
CoOp [43] 15.12 86.53 55.32 37.29 26.20 61.55 75.59 87.00 58.15 59.05 56.18
TPT [25] 17.58 87.02 58.46 40.84 28.33 62.69 74.88 84.49 61.46 60.82 57.66
DiffTPT [8] 17.60 86.89 60.71 40.72 41.04 63.53 79.21 83.40 62.72 62.67 59.85
TDA [14] 17.61 89.70 57.78 43.74 42.11 68.74 77.75 86.18 62.53 64.18 61.03
DPE [37] 19.80 90.83 59.26 50.18 41.67 67.60 77.83 85.97 64.23 61.98 61.93
BPRE (Ours) 22.05 91.24 59.71 52.96 44.74 69.79 77.59 87.35 64.88 63.13 63.34
$\Delta$ +2.25 +0.41 -1.00 +2.78 +2.63 +1.05 -1.62 +1.17 +0.65 -1.05 +1.41
CLIP-ViT-B/16 23.67 93.35 65.48 44.27 42.01 67.44 83.65 88.25 62.59 65.13 63.58
Ensemble 23.22 93.55 66.11 45.04 50.42 66.99 82.86 86.92 65.63 65.16 64.59
CoOp [43] 18.47 93.70 64.51 41.92 46.39 68.71 85.30 89.14 64.15 66.55 63.88
TPT [25] 24.78 94.16 66.87 47.75 42.44 68.98 84.67 87.79 65.50 68.04 65.10
DiffTPT [8] 25.60 92.49 67.01 47.00 43.13 70.10 87.23 88.22 65.74 62.67 65.47
TDA [14] 23.91 94.24 67.28 47.40 58.00 71.42 86.14 88.63 67.62 70.66 67.53
DPE [37] 28.95 94.81 67.31 54.20 55.79 75.07 86.17 91.14 70.07 70.44 69.40
BPRE(Ours) 29.76 95.17 67.73 55.32 55.94 76.86 86.36 92.48 70.24 71.13 70.10
$\Delta$ +0.81 +0.36 +0.42 +1.12 -2.06 +1.79 -0.87 +1.34 +0.17 +0.47 +0.70
Table 2: Cross-Dataset Generalization Analysis. We evaluate top-1 accuracy (%) across methods using both CLIP backbones (ResNet-50 and ViT-B/16). Bold entries denote superior performance.
4.1 Datasets
Our evaluation protocol encompasses two primary benchmarking scenarios: distribution shift robustness and cross-domain generalization. The distribution robustness assessment utilizes ImageNet [5] and its variants (ImageNet-A [13], -V2 [24], -R [12], and -Sketch [30]). For cross-domain evaluation, we employ a diverse suite of 10 recognition datasets, spanning fine-grained categories (FGVCAircraft [20], StanfordCars [16]), texture (DTD [4]), natural objects (Caltech101 [7], Flowers102 [21], Food101 [1], OxfordPets [22]), scenes (SUN397 [32]), satellite imagery (EuroSAT [11]), and action recognition (UCF101 [26]).
4.2 Implementation Details
We adopt PyTorch as the deep learning framework and utilize CLIP with ResNet-50 and ViT-B/16 backbones as the visual encoders. Following TPT, each test image is augmented into 64 views through random resized cropping (63 augmented + 1 original). The Dynamic Prototype Evolution employs a momentum-based update mechanism (momentum=0.9) with an Multi-Dimensional Quality-Aware Reward Module. For the prototype evolution function, we set temperature $\tau=0.01$. The cache mechanism maintains a queue of size 3 per class, with a normalized entropy threshold of 0.1 for prototype updates. During testing, the reward calculator undergoes a 1000-step warmup period with a minimum reward threshold of 0.1, while the prototype manager performs weighted feature averaging every 10 iterations based on computed rewards. All experiments are conducted on either a single NVIDIA RTX 4090 or A100 GPU. For feature space visualization, we utilize t-SNE with perplexity 30 to project the prototype features into 2D space.
4.3 Comparison with State-of-the-art
To validate the effectiveness of our approach, we conduct comprehensive comparisons against several state-of-the-art test-time adaptation methods for CLIP [23]. As a fundamental baseline, we include CLIP’s zero-shot performance using both single prompt (”a photo of a CLASS”) and prompt ensemble strategies. Our comparative analysis encompasses various methodological categories. In the prompt optimization domain, we consider TPT [25], which leverages self-entropy minimization across augmented views, and its advanced variant DiffTPT [8], which incorporates diffusion-based augmentations. For cache-based approaches, we evaluate TDA [14], which employs training-free adapters with positive-negative cache construction, and DMN-ZS [41], which features backpropagation-free dynamic memory aggregation. We further compare against prototype-focused methods, including TPS [27], which efficiently learns per-class prototype shifts, and DPE [37], which implements dual-modal prototype enhancement during testing. To ensure fair evaluation, all baseline results are directly cited from their original publications.
Method Time Acc Gain
CLIP [23] 11 min 59.81 -
TPT [25] $>$ 10 h 60.74 +0.93
DiffTPT [8] $>$ 20 h 60.80 +0.99
TDA [14] 32 min 61.35 +1.54
DPE [37] 2 h 37 min 63.41 +3.60
BPRE(ours) 3 h 04 min 63.77 +3.96
Table 3: Efficiency comparison on ImageNet [5]. We report the testing time, the achieved accuracy, and the performance gains compared to zero-shot CLIP.
Robustness to Distribution Shifts. In Tab.  1, we present comprehensive evaluations on ImageNet and its four challenging out-of-distribution variants. Our BPRE demonstrates consistent improvements over existing approaches across both ResNet-50 and ViT-B/16 backbones. While baseline CLIP [23] shows reasonable in-domain performance, its effectiveness deteriorates under distribution shifts. Though traditional enhancement methods like Ensemble and CoOp [43] offer moderate improvements, their gains remain limited. Among test-time adaptation methods, BPRE achieves state-of-the-art performance across both ResNet-50 and ViT-B/16 architectures. Specifically, with ResNet-50, BPRE surpasses prompt-tuning based approaches (TPT [25] by 3.81% and DiffTPT [8] by 2.36%), cache-based method TDA [14] by 1.49%, and recent methods like DMN-ZS [41] and DPE [37] by 1.10% and 0.26% respectively. Similar performance gains are observed with ViT-B/16, where BPRE outperforms TPT by 4.06%, DiffTPT by 4.22%, TDA by 1.49%, and demonstrates consistent improvements over DMN-ZS (+1.01%) and DPE (+0.57%). These comprehensive results across different architectures validate BPRE’s superior adaptation capability and robustness. These improvements, though modest in absolute terms, are significant given the already strong performance of recent methods. The consistent gains across different architectures and datasets validate the effectiveness of our approach in handling distribution shifts without requiring additional training data.
Cross-Datasets Generalization. Our extensive evaluation in Tab.  2 demonstrates BPRE’s superior performance across 10 diverse datasets using two prominent backbone architectures. With ResNet-50, BPRE achieves significant improvements over the CLIP baseline (+7.52%) and consistently outperforms state-of-the-art methods, including TPT (+5.68%), DiffTPT (+3.49%), TDA (+2.31%), and DPE (+1.41%). The performance gains are particularly notable on challenging fine-grained datasets, with substantial improvements on Aircraft (+2.25%) and DTD (+2.78%). For ViT-B/16, BPRE establishes a new state-of-the-art benchmark with 70.10% average accuracy, demonstrating consistent advantages over leading competitors DPE (+0.70%) and TDA (+2.57%). The method exhibits remarkable domain adaptability, showing significant improvements on specialized datasets like Flower (+1.79%) and Pets (+1.34%). These comprehensive results validate BPRE’s effectiveness as a lightweight test-time adaptation approach that requires neither additional training data nor complex optimization strategies.
Efficiency Comparison. Tab.  3 compares the efficiency of our method with other VLM test-time adaptation approaches on 5,000 ImageNet test samples using a 24GB NVIDIA RTX 4090 GPU. Our model requires only 27 minutes more than the latest DPE [37] model while achieving a 0.36% improvement in prediction accuracy.
4.4 Ablation Studies
Effectiveness of the Multi-Dimensional Quality-Aware Reward Module. To evaluate the contribution of each component in our proposed Multi-Dimensional Quality-Aware Reward Module, we performed an extensive analysis on the ImageNet dataset. As presented in Tab. 4, the removal of the $R_{sim}$, $R_{conf}$, and $R_{div}$ components individually results in a marked degradation in model performance. Specifically, when removing $R_{sim}$, the model’s ability to capture semantic similarities is compromised; without $R_{conf}$, the confidence-based quality assessment is impaired; and the absence of $R_{div}$ reduces the model’s capacity to maintain sample diversity. These findings empirically validate our hypothesis that multi-dimensional evaluation metrics are essential for robust sample quality assessment.
Ablation studies on different values of queue size $M$. Tab. 5 illustrates our systematic investigation of queue size $M$ sensitivity on the ImageNet dataset. Our empirical results demonstrate that $M=3$ yields optimal performance, establishing a critical equilibrium between two competing factors: insufficient intra-class diversity at lower $M$ values and the introduction of detrimental low-confidence samples at higher $M$ values. This finding underscores the importance of appropriate queue size selection in maintaining robust model performance.
# $R_{sim}$ $R_{conf}$ $R_{div}$ Acc.
1 63.48
2 63.56
3 63.59
4 63.52
5 63.61
6 63.65
7 63.64
8 63.77
Table 4: Ablation study of different components of the Multi-Dimensional Quality-Aware Reward Module.
Ablation studies on different values of $R_{min}$. A systematic evaluation of the minimum reward threshold $R_{min}$ was conducted on the ImageNet dataset using a ResNet-50 backbone, as presented in Tab.  6. The experimental results conclusively demonstrate that $R_{min}=0.1$ maximizes model effectiveness. This optimal threshold value plays a crucial role in maintaining the delicate balance between model robustness and discriminative power, highlighting the significance of precise parameter tuning in our proposed framework.
$M$ 1 2 3 4 5
Acc. 63.31 63.52 63.77 63.56 63.72
Table 5: Ablation studies on different values of $M$.
More Sensitivity Analyses of Hyper-Parameters. We conducted a comprehensive hyperparameter sensitivity analysis on the ImageNet dataset using a ResNet-50 backbone, with results visualized in Fig.  5. Fig.  5(a) reveals that the model achieves peak performance when both temperature parameters are set to $\alpha=4$ and $\beta=4$. As illustrated in Fig.  5(b), the optimal balance among reward components is achieved with $\lambda_{sim}=0.6$ and $\lambda_{conf}=0.2$ (with $\lambda_{div}=0.2$ to satisfy the constraint $\lambda_{sim}+\lambda_{conf}+\lambda_{div}=1$).
Values of $R_{min}$ 0.0 0.01 0.05 0.1 0.2 0.3 0.4
ImageNet Acc. 62.76 63.23 63.54 63.77 63.68 63.65 63.59
Table 6: Ablation studies on different values of $R_{min}$.
Figure 5: Hyperparameter analysis of $\alpha$, $\beta$, $\lambda_{sim}$, and $\lambda_{conf}$ was conducted on the ImageNet dataset using a ResNet-50 backbone.
5 Conclusion
In this work, we propose BPRE, a novel framework for test-time adaptation of VLMs that integrates bidirectional prototype evolution with multi-dimensional reward assessment. By incorporating Semantic Similarity and Feature Diversity, BPRE enables more effective feature selection while mitigating text modality biases. The Prototype-Reward Interactive Evolution module further establishes a self-reinforcing cycle, where refined prototypes enhance reward estimation, and improved rewards guide prototype updates, forming a continuous optimization process. Extensive experiments across diverse datasets confirm BPRE’s effectiveness in handling domain shifts while maintaining computational efficiency, making it a promising solution for real-world adaptation scenarios.